OpenStudy (anonymous):

tell whether each equation represents a direct variation. if so, identify the constant of variation: y= 4x +1

OpenStudy (anonymous):

Take a point on this line (2,9) (I randomly picked 2 and put it in for the x and got 9 for the y) Direct Variation says y = kx so 9 = k(2) therefore k = 9/2 Then put in x = 3 and you get y = 13 does 13 = 9/2 * 3? No it doesn't The answer is NO.